OpenStudy (anonymous):

Add. (7x6 + 10x2 – 10) + (3x6 – 6x3 + 4)

OpenStudy (anonymous):

Add. (7x^6 + 10x^2 – 10) + (3x^6 – 6x^3 + 4)

OpenStudy (anonymous):

(42+20-10)+(18-18+4) =52+4 =56

OpenStudy (saifoo.khan):

we have to add like terms.

OpenStudy (saifoo.khan):

(7x^6 + 10x^2 – 10) + (3x^6 – 6x^3 + 4) 7x^6 + 10x^2 - 10 + 3x^6 - 6x^3 + 4 Arrange, 7x^6 + 3x^6 -6x^3 + 10x^2 -10 + 4

OpenStudy (saifoo.khan):

Now, (7+3)x^6 - 6x^3 + 10x^2 - 6 10x^6 -6x^3 + 10x^2 - 6

OpenStudy (anonymous):

oh but that is the answer

OpenStudy (saifoo.khan):

\[10x^6 -6x^3 + 10x^2 - 6 \] ^^^^Asnwer.

OpenStudy (saifoo.khan):

fweaboy, Nope!

OpenStudy (anonymous):

oh ok